# Radiation forces and torques in optics and acoustics

### v1

@@ I.3 About this review
In the past decades, there were a great number of reviews dedicated to: optical forces and manipulation [181, 387, 248, 109, 38, 39, 579, 178, 523, 219, 607, 560, 234, 421, 428, 362, 610, 600], acoustic forces and manipulation [117, 375, 412, 538, 62, 341, 452, 201, 472], acoustofluidic manipulation of bioparticles [211, 328, 184, 271, 585, 318]; and more specifically to: optical pulling forces [187, 183, 332], optical sorting of chiral particles [227, 293], optical binding [208, 180, 109], optical levitation in vacuum [241], optical manipulation of liquid crystal droplets [112] and single molecules [121], plasmonic tweezers [619], optical manipulation with metamaterial structures [489], optomechanics with levitated particles [378], etc.
@@
In this vast literature, which covers most of the specific aspects and applications of optical and acoustic forces, one can see several general omissions. First, very few original papers or reviews have considered optical and acoustic forces on equal footing [543, 179, 2, 538]. Second, radiation torques have received very little attention [413, 121, 452] (mostly in optics) as compared to forces, and have been neglected in the majority of the above reviews. Finally, theoretical introductions in the optical and acoustic forces and torques often lack clear connections with local dynamical properties of the wave fields: the momentum and spin angular momentum densities.
@@
In this review, we aim to fill these gaps and to provide a clear yet comprehensive description of the main optical and acoustic forces and torques. We restrict ourselves to monochromatic but arbitrarily structured wave fields, and consider the fundamental radiation forces and torques in connection with the local momentum and spin properties of the wave fields. Furthermore, we consider nontrivial forces and torques which appear in the cases of complex (anisotropic, chiral, resonant, etc.) particles. Finally, we overview the most important examples of optical and acoustic forces and torques in specific situations: evanescent waves, vortex beams, pulling forces, optical sorting of chiral particles, and acoustofluidic sorting of bioparticles, among others. We hope that our review will provide a clear unified picture of the fundamentals as well as a versatile picture of multiple manifestations and applications of the mechanical action of optical and acoustic fields.
@@
It should be noted that the nomenclature of radiation forces strongly varies across different works and fields. For example, the term “radiation pressure force” can refer to either the gradient force or the force associated with the momentum of the incident wave, while acoustic radiation forces are sometimes referred to as “acoustoforetic”. Throughout this review, we use a unified nomenclature implying: (i) gradient forces associated with gradients of the energy density (intensity) of the incident wave field, (ii) radiation-pressure forces associated with the local momentum (wavevector) of the incident wave, and (iii) recoil forces associated with the momentum carried away by the scattered wave field.
@@ II.1 General features
Throughout this review, we deal with monochromatic wavefields of the form , where is the complex space-variant field amplitude and is the frequency. Quadratic forms of the fields, such as the energy, momentum, force, etc., are calculated assuming time-averaging over one period of oscillations. In this Section, we describe the main dynamical properties of electromagnetic and acoustic (sound) wave fields, which will determine the wave interaction with particles in Section III.
The key quantities of interest are the wave energy, momentum, and spin angular momentum (spin) densities. The wave energy density is typically proportional to the intensity of the wavefield: . However, the definition of wave momentum has produced a number of controversies, both in electromagnetism and acoustics [426, 114, 418, 419, 372, 51, 380]. This is because there are two fundamental types of wave momentum: kinetic and canonical.
@@
The density of the canonical momentum (which is sometimes called ‘pseudomomentum’) is associated with the local phase gradient or local wavevector of the wavefield, Fig. 4(a). It has the form , , which provides a weighted average of the phase gradients of the different components of . It is similar to the probability current in quantum mechanics or the local expectation value of the quantum-mechanical momentum operator , averaged over all field components [76, 78, 86, 87, 103].
@@
In turn, the kinetic momentum density is associated with the energy flux density (e.g., the Poynting vector for electromagnetic fields), which generally has a different form. Remarkably, the relation between the canonical momentum density and kinetic momentum density is described by the universal Belinfante-Rosenfeld relation, first derived in the context of relativistic field theory [71, 518, 76, 86, 87, 119, 103]:
@@
Here, is the spin angular momentum (in what follows, spin) density in the field, which characterizes the local rotation (elliptical polarization) of the vector wave field: . Namely, the time-dependent field traces an ellipse in each point of space, and the spin density is directed along the normal to this ellipse and is proportional to its ellipticity, Fig. 4(b).
@@
Since the canonical and kinetic momentum densities differ by the curl of a vector field, both of these satisfy the stationary continuity equation in a homogeneous lossless medium: , and integral values of these momenta coincide for any localized wave state. Nonetheless, the local values and directions of these momentum densities can differ significantly in inhomogeneous wave fields, and this is crucial for local wave interactions with small particles.
@@
We also note that the total angular momentum density of the wave field consists of the spin and orbital parts [15, 21, 100, 488, 101, 119]: , where the orbital part is determined by the canonical momentum density and the position vector . However, since we will mostly deal with the wave-induced torque with respect to the small-particle center, , it is naturally determined by the spin rather than orbital angular momentum density. The local Belinfante-Rosenfeld relation (1) yields the integral relation for the total angular momentum of a localized wave state expressed via canonical and kinetic momenta: . However, for the local characterization of the wave field, the canonical momentum and spin densities and are the key independent quantities, describing its phase-intensity and polarization features, respectively.
@@
As the vector product “” produces pseudo-vectors, changing their signs under mirror reflections, one can notice that the energy density is a true scalar, the canonical and kinetic momentum densities are true vectors, while the spin and orbital angular momentum densities are pseudo-vectors. One can also construct a pseudo-scalar quadratic form of the wave field: . It characterizes the local chirality of the wave field [551, 530, 98], and plays an important role in interaction with chiral particles not equivalent to their mirror reflections.
@@ II.2 Electromagnetic fields
where and are the permittivity and permeability of the medium, and we write the total energy density as an arithmetic mean of pure electric and magnetic contributions. The kinetic momentum density is described by the well known Poynting vector [279]:
Here we use a Minkowski (rather than Abraham) version of this quantity involving the speed of light in the medium, [426, 380, 51, 89].
The canonical momentum density is given by the arithmetic mean of the general phase-gradient forms for the electric and magnetic fields [76, 86, 87, 89]:
In turn, the electromagnetic spin density is given by the similar electro-magnetic mean of the spin forms [86, 87, 100, 89]:
Using Maxwell’s equations and , one can verify that the quantities (3)–(5) satisfy the Belinfante-Rosenfeld relation (1). Note that the energy, canonical momentum, and spin densities are sums of the electric- and magnetic-field contributions, while the kinetic momentum density (3) mixes the electric and magnetic fields. The factors in the canonical momentum and spin densities (4) and (5) show that these quantities are well defined only for monochromatic fields.
Notably, one can also use alternative definitions of the canonical momentum and spin densities with purely electric (or magnetic) quantities: and . Such quantities also satisfy the Belinfante-Rosenfeld relation (1) with the same kinetic momentum (3) and can be convenient in the problems where magnetic (or electric) properties are inessential [76, 86, 25].
Finally, using Maxwell’s equations transforming curls of the electric/magnetic field into the magnetic/electric field, the chirality density of an electromagnetic field can be written as [98, 97, 123]
This quantity is the helicity density, characterizing the difference between the numbers of right-hand and left-hand circularly polarized photons [6, 551, 123, 86, 17]. For the helicity eigenstates with , (which represent an arbitrary superposition of circularly-polarized plane waves of the same handedness ), the helicity density (6) equals .
| Quantity | Electromagnetism | Acoutics |
@@
| Canonical momentum density | ||
| Spin density | ||
| Chirality (helicity) | 0 |
@@ II.3 Acoustic fields
Acoustic equations (7) and (8), involving the velocity and pressure fields, have the forms similar to their electromagnetic counterparts (2) and (3). However, the canonical momentum and spin densities in sound-wave fields are determined only by the vector velocity field [488, 102, 103]:
@@
Importantly, in addition to the general phase-gradient and polarization-ellipticity explanations, Eqs. (9) and (10) have a clear mechanical interpretation in terms of the microscopic motion of the medium molecules. Although in a plane sound wave the molecules oscillate along one longitudinal direction, in a generic inhomogeneous wave field they follow elliptical trajectories, similar to the polarization ellipses of an optical field. This local motion generates a local mechanical angular momentum with a density given by Eq. (10) [290, 488, 102, 103]. Furthermore, in addition to the linear oscillatory motion, the molecules experience a slow quadratic Stokes drift in the local direction of the wave propagation [520, 113]. The mechanical momentum density corresponding to this drift is precisely the canonical momentum density (9) [103, 93].
Thus, the acoustic canonical momentum and spin densities do not depend explicitly on the scalar pressure field . One can construct macroscopic acoustic field theory with alternative definitions of these densities, with symmetrized contributions from the and fields to the canonical momentum [119], but these definitions contradict the microscopic mechanical calculations. Using the sound wave equations, we note that the kinetic momentum density (8) can be written in a canonical form for the pressure field: . Therefore, to elucidate the analogy with the electromagnetic case and provide symmetric forms to some acoustic-force equations, we will also use notations
Finally, the chirality of the acoustic velocity field vanishes identically, , because this field is curl-less: [101].
@@ III.1 General stress-tensor approach
There is a variety of methods to obtain the radiation forces and torques on matter, such as the Lorentz force approach in optics [181]. However, the most common method, universal to any wave fields, is the stress-tensor approach. When a wave field propagates in a lossless homogeneous isotropic medium, its total energy, momentum, and angular momentum are conserved. This means that the fluxes of these quantities through any closed surface vanish. When the field interacts with a particle (e.g., via absorption or scattering), it can transfer the wave energy, momentum, and angular momentum to the particle. These transfers per unit time are the wave-induced absorption rate, force, and torque on the particle, respectively. They can be calculated as the corresponding fluxes through a closed surface surrounding the particle, Fig. 5 [407, 107, 117]:
Here, is the energy flux density related to the kinetic momentum density:
is the momentum flux density or stress tensor (the flux of a vector quantity is a rank-2 tensor), and is the angular momentum flux density. The stress tensors for monochromatic electromagnetic and acoustic wave fields read [279, 407, 117, 315]:
where is the dyadic product, is the unit tensor, and we note that electromagnetic momentum flux density is equal to the Maxwell stress tensor with opposite sign.
Equations (12)–(14) describe the most general situation with an arbitrary particle and arbitrary wave field. However, the fields in these equations are superpositions of the fields incident and scattered (or radiated) by the particle, Fig. 5:
Since the energy, momentum, and angular momentum fluxes are quadratic functions of the fields, they split into the three corresponding contributions:
[…]
where “mix” contains cross terms with the incident and scattered fields. The “inc” terms in Eq. (16) do not contribute to the integrals (12) because of the conservation laws in the pure incident field. Therefore, the energy absorption, force, and torque on the particle are determined by the mixed and pure scattered contributions, Fig. 5.
Thus, to use Eqs. (12), one first needs to solve the scattering problem to determine the fields scattered by the particle. This problem can be rather complicated in the general case, so that various approximations are often used. In particular, the following scattering regimes can be considered depending on the particle’s characteristic size with respect to the wavelength ( is the wavenumber):
Below we consider various simplified cases, where the scattering problem can be solved analytically. In these cases, the forces and torques can be calculated and expressed in terms of the incident fields.
@@ III.2 General Rayleigh particles
We start with the case of small (Rayleigh) particles with . The scattering is relatively weak and the scattered field can be approximated by point multipoles of the lowest order. In the electromagnetic case, these are electric and magnetic dipoles (oscillating electric/magnetic monopoles do not radiate) [402, 279, 107], whereas for sound waves these are monopole and dipole [574, 83, 467], Fig. 6. These dipole and monopole scattered fields can be characterized by the corresponding dipole and monopole moments: in electromagnetism and in acoustics. We use to denote the electric dipole moment, rather than the commonly accepted , for the sake of unified notations.
Adding the dipole and monopole scattered fields to the incident fields, expanding the incident fields in the first-approximation Taylor series near the particle’s position, and substituting the resulting total fields into Eqs. (12)–(14), one can calculate (see Appendix C) the absorption rate, force, and torque on the particle, which are shown in Table 2 [402, 543, 513]. From now on, we omit the “inc” subscript, and are the incident wave fields taken at the position of the particle, .
The equations in Table 2 are rather general: they only imply smallness of the particle and domination of the dipole and monopole contributions in the scattering field. All other properties of the particle are contained in the dipole and monopole moments. The ‘mixed’ contributions in Table 2 correspond to the minimal-coupling interaction of the point dipoles and monopole with the incident field. In turn, the ‘scattered’ contributions can be interpreted as cross- or self-interaction of the dipoles and monopole. These ‘scattered’ terms have a higher-order smallness in , and will be considered in detail in Subsection III.4.

### v3

@@ I.3 About this review
Over the past decades, there have been numerous reviews dedicated to: optical forces and manipulation [190, 423, 261, 114, 41, 42, 639, 187, 578, 232, 672, 619, 247, 460, 467, 394, 495, 675, 449, 663], acoustic forces and manipulation [123, 410, 450, 593, 65, 370, 211, 520], acoustofluidic manipulation of bioparticles [224, 350, 193, 286, 646, 339]; and more specifically to: optical pulling forces [196, 192, 355], optical lateral forces [542], optical sorting in general [662] and sorting of chiral particles [240, 312], optical binding [221, 189, 114], optical levitation in vacuum [254], optical manipulation of liquid crystal droplets [117] and single molecules [127], plasmonic tweezers [684], optical manipulation with metamaterial structures [541], optomechanics with levitated particles [413], and force-microscopy based on optical trapping [433, 575].
@@
Despite this extensive literature covering most of the specific aspects and applications of optical and acoustic forces, several general gaps are evident. First, only a few original papers and reviews have addressed optical and acoustic forces on equal footing [602, 188, 2, 593]. Second, radiation torques have received comparatively little attention [451, 127, 495] (mostly in optics), and are often neglected in the aforementioned reviews. Finally, theoretical introductions to optical and acoustic forces and torques usually lack clear connections to the local dynamical properties of wave fields: the momentum and spin angular momentum densities.
@@
In this review, we aim to fill these gaps by providing a clear yet comprehensive description of the main optical and acoustic forces and torques. We restrict ourselves to monochromatic but arbitrarily structured wave fields, considering fundamental radiation forces and torques in connection with the local momentum and spin properties of the waves. Additionally, we consider nontrivial forces and torques which appear in the cases of complex (anisotropic, chiral, resonant, etc.) particles. Finally, we overview key examples of optical and acoustic forces and torques in specific settings: evanescent waves, vortex beams, pulling forces, optical sorting of chiral particles, and acoustofluidic sorting of bioparticles, among others. We hope that this review will provide a unified and versatile perspective on both the fundamentals and diverse manifestations of the mechanical action of optical and acoustic fields.
@@
It is worth noting that the nomenclature of radiation forces varies strongly across different studies and disciplines. For example, the term “radiation pressure force” can refer either to the gradient force or to the force associated with the momentum of the incident wave, while acoustic radiation forces are sometimes referred to as “acoustoforetic”. Throughout this review, we adopt a unified nomenclature implying: (i) gradient forces associated with gradients of the energy density (intensity) of the incident wave field, (ii) radiation-pressure forces associated with the local momentum (wavevector) of the incident wave, and (iii) recoil forces associated with the momentum carried away by the scattered wave field.
@@ II.1 General features
In this Section, we outline the main dynamical properties of electromagnetic and sound waves, which will determine the wave interaction with particles in Section III. A more detailed discussion of the momentum and angular momentum of various classical waves can be found in [87].
Throughout this review, we deal with monochromatic wave fields of the form , where is the complex space-variant field amplitude and is the frequency. Quadratic forms of the field, such as energy, momentum, force, etc., are evaluated assuming time-averaging over one oscillation period. The key quantities of interest are the wave energy, momentum, and spin angular momentum (spin) densities. The wave energy density is typically proportional to the field intensity: . However, the definition of wave momentum has long been a subject of debate in both electromagnetism and acoustics [465, 119, 457, 458, 406, 54, 415], primarily because there are two distinct forms: kinetic and canonical momentum.
@@
The density of the canonical momentum (sometimes called ‘pseudomomentum’) is associated with the local phase gradient or local wavevector of the wave field, Fig. 4(a). It can be written as , , which represents a weighted average of the phase gradients of the field components. This form is analogous to the probability current in quantum mechanics or to the local expectation value of the quantum-mechanical momentum operator , averaged over the field components [79, 81, 90, 91, 87].
@@
In turn, the kinetic momentum density is associated with the energy flux density (e.g., the Poynting vector for electromagnetic fields), which generally has a different form. Remarkably, the canonical momentum density and kinetic momentum density are related through the universal Belinfante-Rosenfeld relation, originally derived in relativistic field theory [74, 572, 79, 90, 91, 108, 87]:
@@
Here, is the spin angular momentum (hereafter, spin) density of the field, which characterizes the local rotation (elliptical polarization) of the vector wave field: . Namely, the time-dependent field traces an ellipse at each point of space, and the spin density is directed along the normal to this ellipse and has magnitude proportional to its ellipticity, Fig. 4(b).
@@
Since the canonical and kinetic momentum densities differ by the curl of a vector field, both satisfy the stationary continuity equation in a homogeneous lossless medium: , and their integral values coincide for any localized wave state. However, their local values and directions can differ significantly in inhomogeneous wave fields, and this is crucial for local wave-particle interactions.
@@
The total angular momentum density of the wave field comprises spin and orbital parts [16, 23, 105, 540, 106, 125]: , where the orbital part is determined by the canonical momentum density and the position vector . Since we are primarily concerned with wave-induced torques evaluated relative to the center of a small particle, , such torque is naturally determined by the spin rather than orbital angular momentum density. The local Belinfante-Rosenfeld relation (1) yields the integral identity for the total angular momentum of a localized wave state expressed via canonical and kinetic momenta: . However, for the local characterization of the wave field, the canonical momentum density and spin density remain the two key independent quantities, describing the field’s phase-intensity and polarization features, respectively.
@@
As the cross product “” produces pseudo-vectors, changing their signs under mirror reflections, one can notice that the energy density is a true scalar, the canonical and kinetic momentum densities are true vectors, while the spin and orbital angular momentum densities are pseudo-vectors. One can also construct a pseudo-scalar quadratic form: , which characterizes the local chirality of the field [610, 585, 103]. This quantity plays an important role in interactions with chiral particles that are not equivalent to their mirror reflections.
@@ II.2 Electromagnetic fields
where and are the permittivity and permeability of the medium, and we write the total energy density as the arithmetic mean of the purely electric and purely magnetic contributions. The kinetic momentum density is described by the well-known Poynting vector [296]:
Here we use the Minkowski (rather than Abraham) version of this quantity involving the speed of light in the medium, [465, 415, 54, 93].
Standard electromagnetic field theory yields the canonical momentum and spin densities described by the general forms considered in Section II.1 and involving the electric field [572, 90, 87]:
These quantities are convenient in problems where the matter interacts directly with the wave electric field (e.g., for electric-dipole particles), whereas magnetic coupling is inessential [257, 47, 79, 89, 133, 27].
In more general cases involving both electric and magnetic interactions, it is advantageous to define the canonical momentum and spin densities incorporating both the electric and magnetic fields [79, 129, 90, 91, 105, 11, 93]:
| Quantity | Electromagnetism | Acoustics |
@@
| Canonical momentum density | , | , |
| Spin density | , | |
| Chirality (helicity) |
Using Maxwell’s equations and , one can verify that the quantities (4) and (5), as well as (6) and (5), satisfy the Belinfante-Rosenfeld relation (1) with the kinetic momentum (3). Note that the factors in the canonical momentum and spin densities (4)–(7) indicate that these quantities are well defined only for monochromatic fields.
Finally, by using Maxwell’s equations transforming the curls of the electric/magnetic field into the magnetic/electric field, the chirality density of an electromagnetic field can be expressed as [103, 101, 129, 19]
This is the helicity density, characterizing the difference between the numbers of right-hand and left-hand circularly polarized photons [6, 610, 129, 90, 19, 219]. For helicity eigenstates satisfying , (representing arbitrary superpositions of circularly-polarized plane waves of the same handedness ), the helicity density (8) equals .
@@ II.3 Acoustic fields
Equations (9) and (10), involving the velocity and pressure fields, have forms similar to their electromagnetic counterparts (2) and (3). In turn, the canonical momentum and spin densities in sound waves are determined by the vector velocity field, analogously to Eqs. (4) and (5) [540, 108, 87]:
@@
In addition to the general phase-gradient and polarization-ellipticity explanations, Eqs. (11) and (12) have a clear mechanical interpretation in terms of the microscopic motion of the medium’s molecules. Although in a plane sound wave the molecules oscillate along one longitudinal direction, in a generic inhomogeneous wave field they trace elliptical trajectories, analogous to the polarization ellipses of an optical field. This local motion generates a mechanical angular momentum density given by Eq. (12) [307, 540, 108, 87]. Moreover, in addition to the linear oscillatory motion, the molecules undergo a slow quadratic Stokes drift along the local wave-propagation (phase-gradient) direction [574, 118]. The mechanical momentum density corresponding to this drift is precisely the canonical momentum density (11) [108, 97, 87].
Although canonical momentum and spin densities (11) and (12) do not explicitly depend on the scalar pressure field , one can alternatively define these quantities with contributions from both the and fields [125, 602], analogously to the electric and magnetic contributions in Eqs. (6) and (7). This is helpful in problems involving couplings between matter and both the velocity and pressure fields [602, 123, 410, 593]. Using the sound wave equations, the kinetic momentum density (10) can be written in canonical form for the pressure field:
whereas the spin density associated with the pressure field vanishes identically, , due to the scalar nature of . Finally, the chirality of the acoustic velocity field vanishes as well, , because this field is curl-less: [106].
@@ III.1 General stress-tensor approach
There are various methods to calculate the radiation forces and torques on matter, such as the Lorentz force approach in optics [190] or analysis based on optomechanical work effects [484]. However, the most common method, universal to any wave fields, is the stress-tensor approach. When a wave field propagates in a lossless homogeneous isotropic medium, its total energy, momentum, and angular momentum are conserved. This implies that the fluxes of these quantities through any closed surface vanish. When the field interacts with a particle (e.g., via absorption or scattering), it can transfer energy, momentum, and angular momentum to the particle. The rates of these transfers per unit time correspond to the wave-induced absorption rate, force, and torque on the particle, respectively. They can be calculated as the corresponding fluxes through a closed surface enclosing the particle, Fig. 5 [444, 112, 123]:
Here, is the energy flux density related to the kinetic momentum density as:
is the momentum flux density or stress tensor (the flux of a vector quantity is a rank-2 tensor), and is the angular momentum flux density. For monochromatic electromagnetic and acoustic wave fields, the time-averaged stress tensors take the form [296, 444, 123, 336]:
where is the dyadic product, is the unit tensor, and we note that the electromagnetic momentum flux density equals the Maxwell stress tensor with opposite sign.
Equations (14)–(16) describe the most general case of an arbitrary particle in an arbitrary wave field. However, the fields in these equations are superpositions of the fields incident and scattered (or radiated) by the particle, Fig. 5:
Since the energy, momentum, and angular momentum fluxes are quadratic functions of the fields, each of them can be decomposed into three contributions:
where “mix” denotes the cross terms between the incident and scattered fields. The “inc” terms in Eq. (18) make no contribution to the integrals (14) because of the conservation laws in the unperturbed incident field. Therefore, the energy absorption, force, and torque on the particle are determined by the mixed and pure scattered contributions, Fig. 5.
To apply Eqs. (14), one must first solve the scattering problem to obtain the scattered fields. In general, this problem can be mathematically challenging, so various approximations are often employed. In particular, the following scattering regimes can be distinguished depending on the particle’s characteristic size with respect to the wavelength ( is the wavenumber):
| —Rayleigh approximation | ||||
| —Mie resonant regime | (19) | |||
| —ray approximation |
Below we examine various simplified cases in which the scattering problem can be solved analytically. In these situations, the forces and torques can be expressed directly in terms of the incident fields.
@@ III.2 General Rayleigh particles
We start with the case of small (Rayleigh) particles with . In this regime, the scattering is relatively weak, and the scattered field can be approximated by the lowest-order point multipoles. For electromagnetic waves, these are electric and magnetic dipoles (oscillating electric/magnetic monopoles do not radiate) [439, 296, 112], whereas for sound waves these are monopole and dipole [634, 86, 515], Fig. 6. These scattered fields can be characterized by the corresponding dipole and monopole moments: in electromagnetism and in acoustics. Here we denote the electric dipole moment by , rather than the more conventional , for consistency with a unified notation.
By adding the dipole and monopole scattered fields to the incident fields, expanding the incident fields in the first-order Taylor series about the particle’s position, and substituting the resulting total fields into Eqs. (14)–(16), one can calculate (see Appendix C) the absorption rate, force, and torque on the particle. The resulting expressions are summarized in Table 2 [439, 602, 567]. From now on, we omit the “inc” subscript, and refer to the incident wave fields evaluated at the position of the particle.
The equations in Table 2 are rather general: they only require the particle to be small, with the scattering dominated by the dipole and monopole terms. All other properties of the particle are encoded in its dipole and monopole moments. The ‘mixed’ contributions in Table 2 correspond to the minimal-coupling interaction between the point dipoles/monopole and the incident field. In turn, the ‘scattered’ contributions can be interpreted as cross- or self-interaction of the multipoles. These ‘scattered’ terms are of a higher order in and will be discussed in detail in Section III.4.